# Traces of vanishing Hölder spaces

### v2

@@ 1.2. Basic definitions and main results
Since $|x-y|\geq M-R,$ using condition (5), we can choose $M\gg R$ large enough so that the last term is smaller than $\varepsilon.$ Thus, we have the desired estimate for all possible $z\in[x,y],$ and so (11) gives that the restriction of $\{F,DF,\ldots,D^{m}F\}$ to $E$ belongs to $\dot{\operatorname{VJ}}_{\operatorname{far}}^{m,\omega}(E,V).$
@@
Indeed, this is possible since we are assuming that $V$ is a Banach space, and thus the functions $A_{k}\in(A_{1},\ldots,A_{m})\in\mathcal{A}$ of a jet map Cauchy sequences of $E$ to Cauchy sequences of $\mathcal{L}^{k}(\mathbb{R}^{n},V).$
@@
$D^{\beta}G(x)=\frac{\partial^{|\beta|}G}{\partial x_{1}^{\beta_{1}}\cdots\partial x_{n}^{\beta_{n}}}(x)=D^{|\beta|}G(x)(e_{1},\overset{\beta_{1}}{\ldots},e_{1},\ldots,e_{n},\overset{\beta_{n}}{\ldots},e_{n}),$
@@
where the $A_{\beta}$s are defined as follows. For any multi-index $\beta=(\beta_{1},\dots,\beta_{n})\in(\mathbb{N}\cup\{0\})^{n}$ of order $|\beta|:=\sum_{j}\beta_{j}\leq m,$ and $y\in E,$ we set
$A_{\beta}(y):=A_{|\beta|}(y)(e_{1},\overset{\beta_{1}}{\ldots},e_{1},\ldots,e_{n},\overset{\beta_{n}}{\ldots},e_{n});$ (20)
@@
The second identity in (19) follows from the symmetry and the $k$-linearity of the $A_{k}$s’ and a basic combinatorial argument. Thus, for every multi-index $\alpha\in(\mathbb{N}\cup\{0\})^{n}$ of order $|\alpha|=m,$ one has $D^{\alpha}P_{y}(x)=A_{\alpha}(y),$ for every $x\in\mathbb{R}^{n}$ and $y\in E.$ Now, to show the point $(a)$ it is enough to show that for each multi-index $\alpha$ of order $|\alpha|=m,$ there holds that
@@
We will use property (iii), which tells us that if $x\in Q^{*},$ then $|p_{Q}-y|\leq C|x-y|,$ for some absolute constant $C>0.$ Let $\xi=\xi(x)\in E$ minimize $\operatorname{\mathit{d}}(x,E)=|x-\xi|.$ Together with Lemma 3.2, and the definition of $\operatorname{S},$ we find that
@@
Let us begin with a computation for arbitrary $z\in\mathbb{R}^{n}\setminus E.$ Let $\xi=\xi(z)\in E$ minimize $\operatorname{\mathit{d}}(z,E)=|z-\xi|.$ Then, by the formula (16) and the properties (iii), (v) and (vi) we have
@@
For each cube $Q\in\mathcal{Q}$ so that $x\in Q^{*},$ we have $|p_{Q}-y|\leq C|x-y|$ for an absolute constant $C;$ see (iii). Let $\xi=\xi(x)\in E$ minimize $\operatorname{\mathit{d}}(x,E)=|x-\xi|.$ In this case, although we are letting $|x-y|\to\infty,$ we do not know whether $|p_{Q}-y|$ or $|p_{Q}-\xi|$ are large or not. Therefore, we need to study the behaviour of $|p_{Q}-z|$, for $z\in\{y,\xi\}$ as $|x-y|\to\infty.$ If $|p_{Q}-z|\leq M$ (for a constant $M$ that we will specify in a moment), then we have the bound
$\frac{R(\mathcal{A},z,p_{Q})}{\omega(|x-y|)}\leq\|\mathcal{A}\|_{\dot{\operatorname{J}}^{m,\omega}(E,V)}\frac{\omega(|z-p_{Q}|)}{\omega(|x-y|)}\lesssim\frac{\omega(M)}{\omega(|x-y|)}.$ (27)
On the other hand, if $|p_{Q}-z|\geq M,$ then $\operatorname{L}(|p_{Q}-z|)\leq\operatorname{L}(M)$ and so
$\frac{R(\mathcal{A},z,p_{Q})}{\omega(|x-y|)}\leq\operatorname{L}(|p_{Q}-z|)\frac{\omega(|p_{Q}-z|)}{\omega(|x-y|)}\leq\operatorname{L}(M)\frac{\omega(C|x-y|)}{\omega(|x-y|)}\lesssim\operatorname{L}(M).$ (28)
We have used that if $z\in\{y,\xi\},$ then $|p_{Q}-z|\leq C|x-z|\leq C|x-y|,$ by virtue of (iii). Now, given $\varepsilon>0,$ we choose $M>0$ large enough so that $\operatorname{L}(M)\leq\varepsilon.$ Using $\omega(\infty)=\infty,$ we find $K>M$ such that $\omega(|x-y|)\geq\varepsilon^{-1}\omega(M),$ provided that $|x-y|\geq K.$ Thus the bounds (27) and (28) tell us that regardless of the size of $|p_{Q}-z|,$ we have
@@
From the bound (29) and Lemma 3.2, we obtain
$\frac{\|D^{m}F(x)-D^{m}F(y)\|}{\omega(|x-y|)}\lesssim\sum_{Q^{*}\ni x}\left(\frac{R(\mathcal{A},y,p_{Q})}{\omega(|x-y|)}+\frac{R(\mathcal{A},\xi,p_{Q})}{\omega(|x-y|)}\right)\lesssim_{n}\varepsilon,$
@@
Case 2. Assume $x,y\in\mathbb{R}^{n}\setminus E,$ and $\operatorname{\mathit{d}}([x,y],E)\leq|x-y|.$ Let $p=p(x,y)\in E$ minimize $\operatorname{\mathit{d}}([x,y],E).$ Then, again, $|x-p|,|y-p|\leq 2|x-y|.$ Similarly, we saw already in Case 1 the following: given $\varepsilon>0,$ there exists $M>0$ (independent of $x\in\mathbb{R}^{n}\setminus E$ and $p\in E$) such that: if $|x-p|\geq M,$ then
@@
Let $z\in[x,y]$ and let $\xi=\xi(z)\in E$ minimize $\operatorname{\mathit{d}}(z,E)=|z-\xi|.$ Employing first Lemma 3.2, then (33), and finally property (ii), we derive
@@
By property (13) of $\omega$ and $\operatorname{\mathit{d}}(z,E)\geq|x-y|,$ for all $z\in[x,y],$ we continue bounding,
@@
Case 1. Let $x\in\mathbb{R}^{n}\setminus E,$ $y\in E$. Given $\varepsilon>0,$ let $M>0$ be such that $\mathrm{D}(M)\leq\varepsilon$. We consider $|x|\geq K\gg M$ for a large constant $K$ to be soon determined. Let $\xi=\xi(x)\in E$ minimize $\operatorname{\mathit{d}}(x,E)=|x-\xi|.$ Applying Lemma 3.2 and bearing in mind the property (ii), we find a cube $Q_{x}\in\mathcal{Q}$ with $x\in(Q_{x})^{*}$, and $z\in\{\xi,y\}$ so that

### v3

@@ 1.2. Basic definitions and main results
But observe that, by the triangle inequality and (15), we get $|x-y|\geq M-R\geq M/2.$ In combination with the second inequality of (15), we infer that the last term is bounded above by
$\|F\|_{\dot{C}^{m,\omega}(\mathbb{R}^{n},V)}\frac{\omega(2R)}{\omega(|x-y|)}\leq\|F\|_{\dot{C}^{m,\omega}(\mathbb{R}^{n},V)}\frac{\omega(2R)}{\omega(M/2)}\leq\|F\|_{\dot{C}^{m,\omega}(\mathbb{R}^{n},V)}\varepsilon\lesssim\varepsilon.$
Thus, we have the desired estimate for all possible $z\in[x,y],$ and so (13) gives that the restriction of $\{F,DF,\ldots,D^{m}F\}$ to $E$ belongs to $\dot{\operatorname{VJ}}_{\operatorname{far}}^{m,\omega}(E,V).$
@@
Indeed, this is possible since we are assuming that $V$ is a Banach space, and thus the functions $A_{k}\in(A_{1},\ldots,A_{m})\in\mathcal{A}$ map Cauchy sequences of $E$ to Cauchy sequences of $\mathcal{L}^{k}(\mathbb{R}^{n},V).$
@@
$D^{\beta}G(x)=\frac{\partial^{|\beta|}G}{\partial x_{1}^{\beta_{1}}\cdots\partial x_{n}^{\beta_{n}}}(x)=D^{|\beta|}G(x)\big{(}\overbrace{e_{1},\ldots,e_{1}}^{\beta_{1}\text{ times}},\ldots,\overbrace{e_{n},\ldots,e_{n}}^{\beta_{n}\text{ times}}\big{)},$
@@
where the $A_{\beta}$s are defined as follows. For any multi-index $\beta=(\beta_{1},\dots,\beta_{n})\in(\mathbb{N}\cup\{0\})^{n}$ of order $|\beta|:=\sum_{j=1}^{n}\beta_{j}\leq m,$ and $y\in E,$ we set
$A_{\beta}(y):=A_{|\beta|}(y)(\overbrace{e_{1},\ldots,e_{1}}^{\beta_{1}\text{ times}},\ldots,\overbrace{e_{n},\ldots,e_{n}}^{\beta_{n}\text{ times}});$ (23)
@@
The second identity in (22) follows from the symmetry and the $k$-linearity of the $A_{k}$s’ and a basic combinatorial argument. Thus, for every multi-index $\alpha\in(\mathbb{N}\cup\{0\})^{n}$ of order $|\alpha|=m,$ one has $D^{\alpha}P_{y}(x)=A_{\alpha}(y),$ for every $x\in\mathbb{R}^{n}$ and $y\in E.$ Now, to show part (a) it is enough to show that for each multi-index $\alpha$ of order $|\alpha|=m,$ there holds that
@@
By the property 3.1 in Proposition (iii), if $x\in Q^{*}$ and $u\in E,$ then $|p_{Q}-u|\leq C|x-u|,$ for some absolute constant $C>0.$ Applying this when $u$ is equal to the given $y,$ and also when $u=\xi$ for some $\xi=\xi(x)\in E$ that minimizes $\operatorname{\mathit{d}}(x,E)=|x-\xi|,$ we obtain that
$|p_{Q}-y|\leq C|x-y|,\quad\text{and}\quad|p_{Q}-\xi|\leq C|x-\xi|=C\operatorname{\mathit{d}}(x,E)\leq C|x-y|.$
Together with Lemma 3.2(a), and the definition of $\operatorname{S},$ we find that
@@
Let us begin with a computation for arbitrary $z\in\mathbb{R}^{n}\setminus E.$ Let $\xi=\xi(z)\in E$ minimize $\operatorname{\mathit{d}}(z,E)=|z-\xi|.$ Then, using first Lemma 3.2(b), then the definition of $S$ in (27), and lastly the point (iii) of Proposition 3.1, we get
@@
For each cube $Q\in\mathcal{Q}$ so that $x\in Q^{*},$ we have $|p_{Q}-y|\leq C|x-y|$ for an absolute constant $C;$ see (iii). Let $\xi=\xi(x)\in E$ minimize $\operatorname{\mathit{d}}(x,E)=|x-\xi|.$ In this case, although we are letting $|x-y|\to\infty,$ we do not know whether $|p_{Q}-y|$ or $|p_{Q}-\xi|$ are large or not. Therefore, we need to study the behaviour of $|p_{Q}-z|$, for $z\in\{y,\xi\}$ as $|x-y|\to\infty.$ Given $\varepsilon>0,$ we choose $M>0$ large enough so that $\operatorname{L}(M)\leq\varepsilon.$ Note that such an $M>0$ exists due to the property $\underset{t\to\infty}{\lim}L(t)=0.$ Also, using $\omega(\infty)=\infty,$ we find $K>M$ such that $\omega(t)\geq\varepsilon^{-1}\omega(M),$ provided that $t\geq K.$
Let us assume that $|x-y|\geq K,$ and examine two cases separately. If $|p_{Q}-z|\geq M,$ then $\operatorname{L}(|p_{Q}-z|)\leq\operatorname{L}(M)$ and so
$\frac{R(\mathcal{A},z,p_{Q})}{\omega(|x-y|)}\leq\operatorname{L}(|p_{Q}-z|)\frac{\omega(|p_{Q}-z|)}{\omega(|x-y|)}\leq\operatorname{L}(M)\frac{\omega(C|x-y|)}{\omega(|x-y|)}\lesssim\operatorname{L}(M)\leq\varepsilon.$ (30)
On the other hand, if $|p_{Q}-z|\leq M$, then the assumption $|x-y|>K$ gives
$\frac{R(\mathcal{A},z,p_{Q})}{\omega(|x-y|)}\leq\|\mathcal{A}\|_{\dot{\operatorname{J}}^{m,\omega}(E,V)}\frac{\omega(|z-p_{Q}|)}{\omega(|x-y|)}\lesssim\frac{\omega(M)}{\omega(|x-y|)}\leq\varepsilon.$ (31)
We have used that if $z\in\{y,\xi\},$ then $|p_{Q}-z|\leq C|x-z|\leq C|x-y|,$ by virtue of (iii). Thus the bounds (31) and (30) tell us that regardless of the size of $|p_{Q}-z|,$ we have
@@
From the bound (32), the fact that $D^{m}F(y)=A_{m}(y)$, and Lemma 3.2(a), we obtain
$\frac{\|D^{m}F(x)-D^{m}F(y)\|}{\omega(|x-y|)}=\frac{\|D^{m}F(x)-A_{m}(y)\|}{\omega(|x-y|)}\lesssim\sum_{Q^{*}\ni x}\left(\frac{R(\mathcal{A},y,p_{Q})}{\omega(|x-y|)}+\frac{R(\mathcal{A},\xi,p_{Q})}{\omega(|x-y|)}\right)\lesssim_{n}\varepsilon,$
@@
Case 2. Assume $x,y\in\mathbb{R}^{n}\setminus E,$ and $\operatorname{\mathit{d}}([x,y],E)\leq|x-y|.$ Let $p=p(x,y)\in E$ minimize $\operatorname{\mathit{d}}([x,y],E).$ Then, again, $|x-p|,|y-p|\leq 2|x-y|.$ As in Case 1, given $\varepsilon>0,$ there exists $M>0$ (independent of $x\in\mathbb{R}^{n}\setminus E$ and $p\in E$) such that: if $|x-p|\geq M,$ then
@@
Let $z\in[x,y]$ and let $\xi=\xi(z)\in E$ minimize $\operatorname{\mathit{d}}(z,E)=|z-\xi|.$ Employing first Lemma 3.2(b), then (36), and property (ii) of Proposition 3.1, we derive
@@
By property (16) of $\omega$ and $\operatorname{\mathit{d}}(z,E)\geq|x-y|,$ for all $z\in[x,y],$ we continue,
@@
Case 1. Let $x\in\mathbb{R}^{n}\setminus E,$ $y\in E$. Given $\varepsilon>0,$ let $M>0$ be such that $\mathrm{D}(M)\leq\varepsilon$. We consider $|x|\geq K\gg M$ for a large constant $K$ to be soon determined. Let $\xi=\xi(x)\in E$ minimize $\operatorname{\mathit{d}}(x,E)=|x-\xi|.$ Applying Lemma 3.2(a) and bearing in mind the property (ii), we find a cube $Q_{x}\in\mathcal{Q}$ with $x\in(Q_{x})^{*}$, and $z\in\{\xi,y\}$ so that